# Censoring heavy-tail count distributions for parameter estimation with an application to stable distributions

### v3

@@ Appendix C Proofs
holds. In other words, thanks to continuity of $f_{1}$ and to condition (6), which implies $\lim_{n}p_{*}=p$ almost surely, it suffices to prove
@@
(see, e.g., Marcheselli et al., 2008, page 824). From Glivenko-Cantelli Theorem, $\sup_{p\in]0,1/2]}|\widehat{g}(1-p)-g(1-p)|$ converges to 0 a.s. In particular, if $\lambda^{-1/a}>1/2$, asymptotically $p^{*}=1/2$ a.s.
Moreover, if $\lambda^{-1/a}<1/2$, then $\widehat{g}(1-p_{*})=e^{-1}$ almost surely for large $n$ and
@@
Since $\widehat{g}^{\prime}$ is not a decreasing function
@@
$\lim_{n}\widehat{g}(1-{\lambda}^{-{1\over a}})=g(1-{\lambda}^{-{1\over a}})\quad a.s.,$
from (22) and (23), $p_{*}$ converges a.s. to ${\lambda}^{-{1\over a}}$. (Note that $p_{*}$ converges a.s. to ${1/2}$ when ${\lambda}^{-{1\over a}}=1/2$). Moreover, from (23) it holds
@@
Then, $\widehat{a}$ and $\widehat{\lambda}$ converge almost surely to $a$ and $\lambda$. Moreover, since

### v4

@@ Appendix C Proofs
holds. In other words, thanks to continuity of $f_{1}$ and to condition (6), which implies $\lim_{n}p_{*}=p$ a.s., it suffices to prove
@@
(see e.g., Marcheselli et al., 2008, page 824). Therefore, owing to Glivenko-Cantelli Theorem it follows
$\sup_{p\in]0,1/2]}|\widehat{g}(1-p)-g(1-p)|\to 0,\ \ a.s.$
In particular, if $\lambda^{-1/a}>1/2$, then asymptotically it holds $p^{*}=1/2$ a.s. Moreover, if $\lambda^{-1/a}<1/2$, then $\widehat{g}(1-p_{*})=e^{-1}$ a.s. for large $n$ and
@@
Since $\widehat{g}^{\prime}$ is non-decreasing it follows
@@
$\lim_{n}\widehat{g}(1-{\lambda}^{-{1\over a}})=g(1-{\lambda}^{-{1\over a}})\quad a.s.$
Thus, from (22) and (23), $p_{*}$ converges a.s. to ${\lambda}^{-{1\over a}}$. (Note that $p_{*}$ converges a.s. to ${1/2}$ when ${\lambda}^{-{1\over a}}=1/2$). Moreover, from (23) it holds
@@
Then, $\widehat{a}$ and $\widehat{\lambda}$ converge a.s. to $a$ and $\lambda$. Moreover, since